Mark is 10 years younger than Larry. Larry's age 8 years from now will exceed twice Mark's age 3 years ago by 4 years. Hold old is each now? Please help with step and answer.
thanks

Answer by ankor@dixie-net.com(22740)

Let M = Mark's age now
Let L = Larry's age now
:
Write an equation for each statement:
:
"Mark is 10 years younger than Larry. "
M = L - 10
:
"Larry's age 8 years from now will exceed twice Mark's age 3 years ago by 4 years."
(L + 8) = 2(M - 3) + 4
L + 8 = 2M - 6 + 4
L + 8 = 2M - 2
L = 2M - 2 - 8
L = 2M - 10
:
Replace M with (L-10) from the 1st statement equation
L = 2(L-10) - 10
L = 2L - 20 - 10
L = 2L - 30
+30 = 2L - L
L = 30 is Larry's present age
then
M = L - 10
M = 30 - 10
M = 20 is Marks age
;
:
Check solution in the statement
"Larry's age 8 years from now will exceed twice Mark's age 3 years ago by 4 years."
30 + 8 = 2(20-3) + 4
38 = 2(17) + 4